OpenStudy (nuccioreggie):

Sue used a probability simulator to roll a 6-sided number cube and flip a coin 100 times. The results are shown in the tables below: Number on the Cube Number of Times Rolled 1 18 2 25 3 12 4 32 5 3 6 10 Heads Tails 56 44 Using Sue's simulation, what is the probability of rolling a 2 on the number cube and the coin landing heads up? 81 over 100 56 over 100 1400 over 10000 1100 over 10000 .

OpenStudy (nuccioreggie):

@ganeshie8

OpenStudy (nuccioreggie):

@welshfella

OpenStudy (welshfella):

P(rolling a 2) = 25/100 P (head) = 56/100 multiply these 2 results